If a:b=3:4 and a:(b+c)=2:5.find the value of a:c?
Answer by stanbon(75887)

If a:b=3:4 and a:(b+c)=2:5.find the value of a:c?
--------
If a:b=3:4 then a/b=3/4 and 4a=3b
If a:(b+c)=2:5 then a/(b+c)=2/5 and 5a=2b+2c
---------
Comment: To find a:c we need an equation that relates a to c----no b's.
Substitute: If 4a=3b then b=(4/3)a
If 5a=2b+2c then 5a=2(4/3)a+2c
5a=(8/3)a+2c
(15/3)a=(8/3)a+2c
(7/3)a=2c
7a=6c
Therefore a:c=6:7
Cheers,
Stan H.